what would the answer be for
given f (x) 5^2-2x+1 find f (-2)

Answer by stanbon(75887)

given f(x)= 5^2-2x+1
find f(-2) = 5^2 - 2*-2 +1 = 25 +4 +1 = 30
=================
Cheers,
Stan H.